# Simplifying generic smooth maps to the $2$–sphere and to the plane

### v1

@@ Abstract.
We study how to construct explicit deformations of generic smooth maps from closed $n$–dimensional manifolds $M$ with $n\geq 4$ to the $2$–sphere $S^{2}$ and show that every smooth map $M\to S^{2}$ is homotopic to a $C^{\infty}$ stable map with at most one cusp point and with only folds of the middle absolute index. Furthermore, if $n$ is even, such a $C^{\infty}$ stable map can be so constructed that the restriction to the singular point set is a topological embedding. As a corollary, we show that for $n\geq 4$ even, there always exists a $C^{\infty}$ stable map $M\to\mathbb{R}^{2}$ with at most one cusp point such that the restriction to the singular point set is a topological embedding. As another corollary, we give a new proof to the existence of an open book structure on odd dimensional manifolds which extends a given one on the boundary, originally due to Quinn. Finally, using the open book structure thus constructed, we show that $k$–connected $n$–dimensional manifolds always admit a fold map into $\mathbb{R}^{2}$ without folds of absolute indices $i$ with $1\leq i\leq k$, for $n\geq 7$ odd and $21\leq k<(n-3)/$ 2.
@@ 1. Introduction
For an arbitrary smooth closed manifold $M$ of dimension $n\geq 1$, we have a Morse function $f:M\to\mathbb{R}$. It is known that by modifying $f$ by homotopy, we can arrange so that the critical values are in the order of indices, i.e. for every pair $p,q$ of critical points of indices $i(p),i(q)$, respectively, with $i(p)<i(q)$, we have $f(p)<f(q)$. Furthermore, Smale [20] showed that if an $n$–dimensional closed manifold $M$ with $n\geq 6$ is $k$–connected with $k\geq 1$, then there exists a Morse function $f:M\to\mathbb{R}$ without critical points of indices $i$ with $1\leq i\leq k$ and $n-k\leq i\leq n-1$.
So, the following problem naturally arises.
@@
How about generic maps (or $C^{\infty}$ stable maps) to manifolds of dimensions $\geq 2$? The singular values can be nicely arranged with respect to the indices? If the manifold is highly connected, then the singular points of certain indices can be eliminated?
@@
We note that such a problem has been considered in [5], for example. Recall that for a $C^{\infty}$ stable map $f$ of an $n$–dimensional manifold $M$, $n\geq 2$, into a manifold of dimension $2$, we have only folds and cusps as its singularities, and each of them has an absolute index (for details, see §2), which is a concept similar to that for a nondegenerate critical point of a smooth function. Furthermore, the singular point set $S(f)$ is a closed $1$–dimensional submanifold of $M$.
@@
In the present paper, we consider the above problem for generic smooth maps to $S^{2}$ and to $\mathbb{R}^{2}$. Our first result is Theorem 4.3: for a closed connected $n$–dimensional manifold $M$ with $n\geq 3$, there always exists a $C^{\infty}$ stable map $f:M\to S^{2}$ such that for $n$ odd, $f$ has only folds of absolute index $(n-1)/2$ and has no cusps and that for $n$ even, $f$ has only folds of absolute index $(n-2)/2$ and at most one cusp, whose absolute index is equal to $(n-2)/2$. Furthermore, when $n$ is even, we can choose $f$ such that the restriction of $f$ to the singular point set $S(f)$ is a topological embedding. In the case of $n$ even, using the latter result, we will show, as Theorem 5.1, that there always exists a $C^{\infty}$ stable map $f:M\to\mathbb{R}^{2}$ with at most one cusp point such that $f|_{S(f)}$ is a topological embedding.
@@
In the case of $n$ odd, we will also show that we can construct a $C^{\infty}$ stable map $f:M\to\mathbb{R}^{2}$ without cusps such that the singular value set winds around the origin. This leads to a new proof to the existence of an open book structure on odd dimensional smooth manifolds (see Theorem 6.3), which is originally due to Winkelnkemper [25], Tamura [22], Lawson [10] and Quinn [14]. In fact, we will show that any open book structure on the boundary $\partial M$ of a compact odd dimensional manifold $M$ extends through $M$. The open book that we construct has a very nice property with respect to radial Morse function, and this observation will be used to prove Theorems 7.2 and 7.6: if $M$ is a closed $k$–connected $n$–dimensional manifold with $n\geq 7$ odd and $21\leq k\leq(n-5)/$ 2, then there exists a $C^{\infty}$ stable map $f:M\to\mathbb{R}^{2}$ without cusp points such that $f$ has no fold points of absolute index $i$ with $1\leq i\leq k$. This gives a good answer to the second half of Problem 1.1.
@@
The paper is organized as follows. In §2, we review important properties of generic maps of manifolds of dimension $n\geq 2$ into $2$–dimensional manifolds. In particular, we give some lists of changes of the singular value set for a generic $1$–parameter family of such maps, which will play an essential role in this paper. In fact, in the case of manifolds $M$ of dimension $n=4$, a similar set of lists has been given in [2] (see also [1]). In §3, we introduce the notion of moves, which change the singular value set of a generic smooth map. We rigorously formulate the notion of an always-realizable move, which was originally introduced in [2], as a formal change of the singular value set that is always realized by a generic $1$–parameter family of smooth maps.
@@
In §4, using such always-realizable moves, we try to arrange the singular value set of generic maps into $S^{2}$. We first arrange them so that they wind around a certain point in the target, and then we arrange them in the order of the absolute index. These explicit procedures using always-realizable moves lead to the proof of Theorem 4.3. In §5, using such a generic smooth map into $S^{2}$, in the even dimensional case, we construct a generic smooth map into $\mathbb{R}^{2}$ and prove Theorem 5.1. The idea of the proof of these theorems are basically similar to that exploited in dimension $4$ in [2]. However, in dimension $4$, one can use Lefschetz singularities (see Remark 4.12). In our case of arbitrary dimensions, we will give new techniques that do not involve Lefschetz singularities. This is one of the new contributions of this paper.
@@
In §6, for a compact manifold $M$ of an odd dimension with an open book structure being given on the boundary, we construct a smooth generic map into $\mathbb{R}^{2}$ which is compatible to the open book structure on the boundary, and, using techniques employed in §4, we arrange the singular value set so that it winds around the origin. Then, we prove the existence of an open book structure on $M$ which extends the given one on $\partial M$. In §7, we study the connectivity of the page and the binding of the open book structure thus constructed, and using the Cerf theory techniques developed in [8], we eliminate fold loci of certain absolute indices.
@@
Throughout the paper, all manifolds and maps are smooth of class $C^{\infty}$ unless otherwise specified. For a smooth map $f:M\to N$ between smooth manifolds, $S(f)(\subset M)$ denotes the set of singular points of $f$. For a real number $x$, $\lfloor x\rfloor$ denotes the greatest integer not exceeding $x$.
@@ 2. Generic maps into surfaces
with respect to appropriate coordinates around $p$ and $f(p)$ is called a fold singularity (or a fold), where the integer $i$ is called the index with respect to the $(-Y)$–direction. When we choose the $Y$–coordinate so that $0\leq i\leq\left\lfloor(n-1)/2\right\rfloor$ holds, $i$ is called the absolute index. [^1] We warn the reader that this convention is different from that adopted in [11, p. 273], [26], etc.
@@
is called a cusp singularity (or a cusp), where $i$ is called the index with respect to the $(-Y)$–direction. When we choose the $Y$–coordinate so that $0\leq i\leq\left\lfloor(n-2)/2\right\rfloor$ holds, $i$ is called the absolute index.
@@
For the image of a cusp point, we can also consider a “normal orientation”: this is considered to be normal to the two branches of curves emanating from the point. The indices around the image of a cusp are always as depicted in Fig. 2 for some integer $i$ with respect to appropriate normal orientations as in the figure.
@@
Figure 2. Normal orientations and indices around the image of a cusp
@@ 3. Diagrams and their moves
Furthermore, two normally oriented indexed diagrams are isotopic if there exists an ambient isotopy of $N$ which takes one to the other including the normal orientations and the indices. We often use the terminology “diagram” in order to refer to its isotopy class.
@@
    $\mathcal{D}$ coincides with $\mathcal{D}^{\prime}$ outside of $\Delta$, including the normal orientations and the indices.
@@
    there exists a disk $\Delta_{f}$ in $N$ such that for the diagram $(N,f(S(f)))$, $\partial\Delta_{f}$ intersects $f(S(f))$ outside of the images of the cusps and double points transversely and that $\Delta_{f}\cap f(S(f))$ is isotopic to $\Delta\cap\mathcal{D}$ under a certain identification of $\Delta_{f}$ with $\Delta$.
Suppose that there exists a smooth $1$–parameter family $f_{t}:M\to N$, $t\in[0,1]$, of smooth maps, with $f_{0}=f$, which modifies the singular point set image $f(S(f))$ only in $\mathop{\mathrm{Int}}\nolimits{\Delta_{f}}$, such that $\Delta_{f}\cap f_{1}(S(f_{1}))$ is isotopic to $\Delta\cap\mathcal{D}^{\prime}$, that $f_{t}^{-1}(\Delta_{f})=f^{-1}(\Delta_{f})$ for all $t$, that $f_{t}=f$ outside of the inverse images of $\Delta_{f}$ for all $t$, and that except for finitely many $t$’s, $f_{t}$ is a stable map. Then, we say that the move $\Delta\cap\mathcal{D}\longrightarrow\Delta\cap\mathcal{D}^{\prime}$ is realized for $f$ (or is realizable). If a move is realized for every $f$ with property ()($\sharp$)( ♯ ) above, then we say that the move is always-realizable.
@@
Some of the following moves are always-realizable, while others are not. The reader should refer to the arguments given in [2], where the case of $n=4$ is thoroughly explained. The arguments for general dimension $n$ are basically similar, although there are some subtle issues for higher dimensions [8].
Fig. 4 gives the first set of moves that we will use later. Note that outside of the small disks depicted in the figure, these moves do not change the diagrams. These are called moves of type I: note that each move corresponds to a generic $1$–parameter family of smooth maps that changes the map only in a neighborhood of a point in the source manifold. The reader is referred to Chapter V and Chapter VII, Proposition 1 of [8] for $n\geq 6$.
@@
Note that the simple arrows are always-realizable, while the arrows with a slanted line are not. “Cusp Merge”, which is depicted by a dotted arrow, is “always-realizable” as long as the fibers over the region indicated by ()($\ast$)( ∗ ) of a given stable map are connected.
The moves as depicted in Fig. 5 are called moves of type II. Note that these moves correspond to generic $1$–parameter families of smooth maps that change the maps in neighborhoods of two points in the source manifold. For these moves and the following, the reader is referred to Chapter V and Chapter I, §7 of [8].
@@
The moves as depicted in Fig. 7 are called cusp–fold crossing moves (refer to [8, Chapter V, (0.3)]).
@@
Figure 8. Exchange move and criss-cross move
@@
Figure 9. Moves for the exchange move
A more detailed discussion about the exchange move can be found in [2, Proposition 3.2] for $n=4$ and $i=1$. The reader is also referred to [8, Chapter V, §4].
@@ 4. Stable maps into $S^{2}$
Recall that a generic map $f:M\to N$ is simple if it has no cusps and each connected component of every fiber contains at most one singular point (see [12]). In the literature, some authors use the terminology “strongly simple” instead of “image simple”.
@@
The above theorem holds also for $n=2$. See [27, Theorem 1.4], for example.
@@
Let us assume that $f:M\to S^{2}$ is a stable map with at most one cusp such that the minimum of the absolute indices of folds is equal to $i_{0}$ with $0<i_{0}<\lfloor(n-1)/2\rfloor$. We successively eliminate fold singularities (by homotopy) in the order of absolute index, without changing the number of cusps, as follows.
@@
Note that when $f$ has exactly one cusp point of index $(n-2)/2$, if $f$ is outward-directed, then the cusp is also “outward” (see Fig. 3 (3)).
@@
Let $f:M\to S^{2}$ be a stable map with at most one cusp point such that the minimum of the absolute index of folds is equal to $i_{0}$ with $0<i_{0}<\lfloor(n-1)/2\rfloor$. Then, by using always-realizable moves we can make the fold image of absolute index $i_{0}$ outward-directed in such a way that the other folds have absolute index strictly greater than $i_{0}$, that their images are situated in the innermost region and that the resulting stable map has at most one cusp.
@@
Let $D\subset S^{2}$ be a $2$–disk containing $f(S(f))$. We can also take an annulus $A\cong[0,1]\times S^{1}$ in $D$ containing $f(S(f))$ in its interior. We use almost the same argument as in the proof of [2, Theorem 4.1] in order to arrange so that the fold image of absolute index $i_{0}$ is contained in $A$, and is outward-directed, and that the other fold images are situated in the inner region of $D\setminus A$. The points different from the situation in [2] that we need to take care of are as follows.
@@
Let $f:M\to S^{2}$ be a stable map with at most one cusp point such that the minimum of the absolute index of folds is equal to $i_{0}$ with $0<i_{0}<\lfloor(n-1)/2\rfloor$. We assume that $M$ is connected. Then, by using always-realizable moves we can eliminate the folds of absolute index $i_{0}$ in such a way that the other folds have absolute index strictly greater than $i_{0}$ and that the resulting stable map has at most one cusp.
@@
We further follow the procedure as described in [2, Figs. 30 and 31]. If the top strand, denoted by $b$, does not intersect the other strands, then we can isotope it into a small open disk neighborhood of the north pole $NP$. Then, we can use the move as described in Fig. 10 of the present paper so as to turn it to a fold image of absolute index $i_{0}+1$ which is an embedded small circle and is directed outward. If the top strand does intersect the others, then we perform the moves as described in [2, Fig. 33], by using the move as in Fig. 11 of the present paper instead of the move as described in [2, Fig. 34]. Note that by this move, we again create an embedded fold image circle of index $i_{0}+1$, which can be isotoped to a neighborhood of the south pole $SP$. In order to reach the second stage of [2, Fig. 33], we need to swing the relevant fold image of index $i_{0}$ over the region around $SP$: this is possible, since all the fold images in this region have absolute index strictly greater than $i_{0}$.
Then, we reach the final stage of [2, Fig. 35]. Finally, we apply the move as in Fig. 10 of the present paper to turn the index $i_{0}$ circle to one of index $i_{0}+1$.
@@
In our present case, there may appear some small embedded fold image circles of index $i_{0}+1$ normally oriented outward (see Figure 11). Note that if we reverse the normal orientation, then the index turns into $i_{0}$, the absolute index, since $n$ is even and $i_{0}=(n-2)/2$. We used push moves under the presence of Lefschetz critical points (see [2, Fig. 2]). In our present case, we can instead use the moves as depicted in Fig. 12.
@@
Then, we reach the situation as described in the leftmost figure of [2, Fig. 36], where the absolute indices are all equal to $i_{0}$ and each component is normally oriented inward. Furthermore, one of the components may contain a cusp image.
This completes the proof. ∎

### v2

@@ Abstract.
We study how to construct explicit deformations of generic smooth maps from closed $n$–dimensional manifolds $M$ with $n\geq 2$ to the $2$–sphere $S^{2}$ and show that every smooth map $M\to S^{2}$ is homotopic to a $C^{\infty}$ stable map with at most one cusp point and with only folds of the middle absolute index. Furthermore, if $n$ is even, such a $C^{\infty}$ stable map can be so constructed that the restriction to the singular point set is a topological embedding. As a corollary, we show that for $n\geq 2$ even, there always exists a $C^{\infty}$ stable map $M\to\mathbb{R}^{2}$ with at most one cusp point such that the restriction to the singular point set is a topological embedding. As another corollary, we give a new proof to the existence of an open book structure on odd dimensional manifolds which extends a given one on the boundary, originally due to Quinn. Finally, using the open book structure thus constructed, we show that $k$–connected $n$–dimensional manifolds always admit a fold map into $\mathbb{R}^{2}$ without folds of absolute indices $i$ with $1\leq i\leq k$, for $n\geq 7$ odd and $21\leq k\leq(n-5)/$ 2.
Key words and phrases:
Generic map, fold, cusp, stable map, absolute index, open book structure
2020 Mathematics Subject Classification:
Primary 57R45; Secondary 58K30, 57R70, 58K05.
@@ 1. Introduction
For an arbitrary smooth closed manifold $M$ of dimension $n\geq 1$, we have a Morse function $M\to\mathbb{R}$. It is known that by modifying such a function by homotopy if necessary, we can arrange so that the critical values are in the order of indices, i.e. for every pair $p,q$ of critical points of indices $i(p),i(q)$, respectively, with $i(p)<i(q)$, we have $f(p)<f(q)$, where $f:M\to\mathbb{R}$ denotes the modified Morse function.
Furthermore, Smale [27] showed that if an $n$–dimensional closed manifold $M$ with $n\geq 6$ is $k$–connected with $k\geq 1$, then there exists a Morse function $f:M\to\mathbb{R}$ without critical points of indices $i$ with $1\leq i\leq k$ and $n-k\leq i\leq n-1$. This was a very important result in differentiable topology, which showed a strong geometric consequence of an algebraic topological property for manifolds, and which led to the solution to the Poincaré Conjecture in high dimensions.
So, the following problem naturally arises. For the definitions and properties of generic maps or $C^{\infty}$ stable maps into manifolds of dimension $2$, see §2.
@@
How about generic maps or $C^{\infty}$ stable maps to manifolds of dimensions $\geq 2$? The singular values can be nicely arranged with respect to the indices? If the manifold is highly connected, then the singular points of certain indices can be eliminated?
@@
We note that such a problem has been considered in [9], for example. Recall that for a $C^{\infty}$ stable map $f$ of an $n$–dimensional manifold $M$, $n\geq 2$, into a manifold of dimension $2$, we have only folds and cusps as its singularities, and each of them has an absolute index (for details, see §2), which is a concept similar to that for a nondegenerate critical point of a smooth function. Furthermore, the singular point set is a closed $1$–dimensional submanifold of $M$.
@@
In the present paper, we consider the above problem for generic smooth maps to $S^{2}$ and to $\mathbb{R}^{2}$. Our first result is Theorem 4.3 (see also Remark 4.4): for a closed connected $n$–dimensional manifold $M$ with $n\geq 2$, there always exists a $C^{\infty}$ stable map $f:M\to S^{2}$ such that for $n$ odd, $f$ has only folds of absolute index $(n-1)/2$ and has no cusps and that for $n$ even, $f$ has only folds of absolute index $(n-2)/2$ and at most one cusp, whose absolute index is equal to $(n-2)/2$. Furthermore, when $n$ is even, we can choose $f$ such that the restriction of $f$ to the singular point set $S(f)$ is a topological embedding. In the case of $n$ even, using the latter result, we will show, as Theorem 5.1 (see also Remark 5.2), that there always exists a $C^{\infty}$ stable map $f:M\to\mathbb{R}^{2}$ with at most one cusp point such that $f|_{S(f)}$ is a topological embedding.
@@
In the case of $n$ odd, we will also show that we can construct a $C^{\infty}$ stable map $f:M\to\mathbb{R}^{2}$ without cusps such that the singular value set winds “monotonically” (with respect to the polar angle) around the origin. This leads to a new proof to the existence of an open book structure on odd dimensional smooth manifolds (see Theorem 6.3), which is originally due to Winkelnkemper [32], Tamura [29], Lawson [14] and Quinn [19]. In fact, we will show that any open book structure on the boundary $\partial M$ of a compact odd dimensional manifold $M$ extends through $M$. The open book structure that we construct has a very nice property with respect to a radial Morse function, and this observation will be used to prove Theorems 7.2 and 7.6: if $M$ is a closed $k$–connected $n$–dimensional manifold with $n\geq 7$ odd and $21\leq k\leq(n-5)/$ 2, then there exists a $C^{\infty}$ stable map $f:M\to\mathbb{R}^{2}$ without cusp points such that $f$ has no fold points of absolute index $i$ with $1\leq i\leq k$. This gives a reasonable answer to the second half of Problem 1.1. It is surprizing to see such a generalization of the above-mentioned result due to Smale, which led to the solution to the higher dimensional Poincaré Conjecture, to generic maps (or $C^{\infty}$ stable maps) into the plane, although they are restricted to odd dimensions, since now the singular points are not discrete and constitute positive dimensional submanifolds. The main ingredient for our construction of deformations of generic maps is the so-called Cerf Theory [4, 12].
@@
The paper is organized as follows. In §2, we review important properties of generic maps of manifolds of dimension $n\geq 2$ into $2$–dimensional manifolds. In particular, we give some lists of changes of the singular value set for a generic $1$–parameter family of such maps, which will play an essential role in this paper. In fact, in the case of manifolds $M$ of dimension $n=4$, a similar set of lists has been given in [3] (see also [2]).
In §3, we introduce the notion of moves, which change the singular value set of a generic smooth map. We rigorously formulate the notion of an always-realizable move, which was originally introduced in [3], as a formal change of the singular value set that is always realized by a generic $1$–parameter family of smooth maps.
@@
In §4, using such always-realizable moves, we start to arrange the singular value set of generic maps into $S^{2}$. We first arrange them so that they wind monotonically around a certain point in the target, and then we arrange them in the order of the absolute index. These explicit procedures using always-realizable moves lead to the proof of Theorem 4.3.
In §5, using such a generic smooth map into $S^{2}$, in the even dimensional case, we construct a generic smooth map into $\mathbb{R}^{2}$ and prove Theorem 5.1. The idea of the proof of these theorems are basically similar to that exploited in dimension $4$ in [3]. However, in dimension $4$, one can use Lefschetz singularities (see Remark 4.12). In our case of arbitrary dimensions, we will give new techniques that do not involve Lefschetz singularities. This is one of the new contributions of this paper.
@@
In §6, for a compact manifold $M$ of odd dimension with an open book structure being given on the boundary, we construct a smooth generic map into $\mathbb{R}^{2}$ which is compatible with the open book structure on the boundary, and, using techniques employed in §4, we arrange the singular value set so that it winds monotonically around the origin. Then, we prove the existence of an open book structure on $M$ which extends the given one on $\partial M$.
In §7, we study the connectivity of the page and the binding of the open book structure thus constructed, and using the Cerf theory techniques developed in [12], we eliminate fold loci of certain absolute indices.
Finally, in §8, we consider the topological position of the singular point set of a generic map, instead of its image. Such study was conducted in the author’s previous paper [21]; however, in some cases, the result was not correctly stated. In this section, we give a correction to the result and its proof.
@@
Throughout the paper, all manifolds and maps are smooth of class $C^{\infty}$ unless otherwise specified. The symbol $\cong$” denotes a diffeomorphism between manifolds or an appropriate isomorphism between algebraic objects. All homology groups are with integer coefficients unless otherwise specified. For a smooth map $f:M\to N$ between smooth manifolds, $S(f)(\subset M)$ denotes the set of singular points of $f$. For a real number $x$, $\lfloor x\rfloor$ denotes the greatest integer not exceeding $x$.
@@ 2. Generic maps into surfaces
with respect to appropriate coordinates around $p$ and $f(p)$ is called a fold singularity (or a fold), where the integer $i$ is called the index with respect to the $(-Y)$–direction. When we choose the $Y$–coordinate so that $0\leq i\leq\left\lfloor(n-1)/2\right\rfloor$ holds, $i$ is called the absolute index. [^1] We warn the reader that this convention is different from that adopted in [15, p. 273], [33], etc. Note that the absolute index of a fold is well-defined and does not depend on the choice of the local coordinates.
@@
is called a cusp singularity (or a cusp), where $i$ is called the index with respect to the $(-Y)$–direction. When we choose the $Y$–coordinate so that $0\leq i\leq\left\lfloor(n-2)/2\right\rfloor$ holds, $i$ is called the absolute index, which is well-defined.
@@
For the image of a cusp point, we can also consider a “normal orientation”: this is considered to be normal to the two branches of curves emanating from the point. The indices of the folds around the image of a cusp are always as depicted in Fig. 2 for some integer $i$ with respect to appropriate normal orientations as in the figure.
@@
Figure 2. Indices and normal orientations around the image of a cusp
@@ 3. Diagrams and their moves
Furthermore, two normally oriented indexed diagrams are isotopic if there exists an ambient isotopy of $N$ which takes one to the other preserving the normal orientations and the indices. We often use the terminology “diagram” in order to refer to its isotopy class.
@@
    $\mathcal{D}$ coincides with $\mathcal{D}^{\prime}$ outside of $\Delta$, together with the normal orientations and the indices.
@@
    there exists a disk $\Delta_{f}$ in $N$ such that for the diagram $(N,f(S(f)))$, $\partial\Delta_{f}$ intersects $f(S(f))$ outside of the images of the cusps and double points transversely and that under a certain identification of $\Delta_{f}$ with $\Delta$, $\Delta_{f}\cap f(S(f))$ is isotopic to $\Delta\cap\mathcal{D}$, relative to the boundary $\partial\Delta_{f}\cong\partial\Delta$.
Suppose that there exists a smooth $1$–parameter family $f_{t}:M\to N$, $t\in[0,1]$, of smooth maps, with $f_{0}=f$, which modifies the singular point set image $f(S(f))$ only in $\mathop{\mathrm{Int}}\nolimits{\Delta_{f}}$, such that $\Delta_{f}\cap f_{1}(S(f_{1}))$ is isotopic to $\Delta\cap\mathcal{D}^{\prime}$ relative to the boundary $\partial\Delta_{f}\cong\partial\Delta$, that $f_{t}^{-1}(\Delta_{f})=f^{-1}(\Delta_{f})$ for all $t$, that $f_{t}=f$ outside of the inverse images of $\Delta_{f}$ for all $t$, and that except for finitely many $t$’s, $f_{t}$ is a stable map. Then, we say that the move $\Delta\cap\mathcal{D}\longrightarrow\Delta\cap\mathcal{D}^{\prime}$ is realized for $f$ (or is realizable). If a move is realized for every $f$ with property ()($\sharp$)( ♯ ) above, then we say that the move is always-realizable.
@@
Some of the following moves are always-realizable, while others are not. The reader should refer to the arguments given in [3], where the case of $n=4$ is thoroughly explained. The arguments for general dimension $n$ are basically similar, although there are some subtle issues for higher dimensions. For details, the reader is referred to [4, 5, 12].
Fig. 4 gives the first set of moves that we will use later. Note that outside of the small disks depicted in the figure, these moves do not change the diagrams. These are called moves of type I: note that each move corresponds to a generic $1$–parameter family of smooth maps that changes the map only in a neighborhood of a point in the source manifold. The reader is referred to [12, Part I, Chapter V and Chapter VII, Proposition 1] for $n\geq 6$.
@@
Note that the simple arrows are always-realizable, while the arrows with a slanted line are not. “Cusp Merge”, which is depicted by a dotted arrow, is “always-realizable” as long as the fibers over the region indicated by ()($\ast$)( ∗ ) of a given stable map are connected. Furthermore, “Unflip” is ‘always-realizable” as long as the index satisfies the assumption $i=0$ or $22\leq i\leq n-$ 2 (see [12, Part I, Chapter V, Proposition 1.4]).
The moves as depicted in Fig. 5 are called moves of type II. Note that these moves correspond to generic $1$–parameter families of smooth maps that change the maps in neighborhoods of two points in the source manifold. For these moves and the following, the reader is referred to [12, Part I, Chapter V and Chapter I, §§7 and 8] and [4, Chapitre IV, §2.2, Proposition 2].
@@
The important feature of the moves of type II (and those of type III that follows) is that fold image of smaller index can be pushed to the direction of the normal orientation.
@@
The moves as depicted in Fig. 7 are called cusp–fold crossing moves (refer to [12, Part I, Chapter V, (0.3)] and [4, Chapitre IV, Proposition 4]).
@@
Figure 8. Exchange move and criss-cross move, which are valid for $i\geq 1$.
@@
Figure 9. Moves for the exchange move, $i\geq 1$.
A more detailed discussion about the exchange move can be found in [3, Proposition 3.2] for $n=4$ and $i=1$. For higher dimensions, the case of $i=1$ can be proved by the same argument as in [3]. For $i\geq 2$, we can apply flips, cusp merges, a move of type III, and an unflip safely, since in our situation they are all “always-realizable.”
@@ 4. Stable maps into $S^{2}$
Recall that a generic map $f:M\to N$ is simple if it has no cusps and each connected component of every fiber contains at most one singular point (see [16]). In the literature, some authors use the terminology “strongly simple” instead of “image simple”. Note that a generic map which is image simple and has no cusps is simple.
@@
Theorem 4.3 holds also for $n=2$. See [34, Theorem 1.4], for example.
@@
Let us assume that $f:M\to S^{2}$ is a stable map with at most one cusp such that the minimum $i_{0}$ of the absolute indices of folds satisfies $0<i_{0}<\lfloor(n-1)/2\rfloor$. We successively eliminate fold singularities (by homotopy) in the order of absolute index, without changing the number of cusps, as follows.
@@
Note that when $f$ has exactly one cusp point of index $(n-2)/2$, if the fold image of absolute index $(n-2)/2$ of $f$ is outward-directed, then the cusp image also points “outward” (see Fig. 3 (3)).
@@
Let $f:M\to S^{2}$ be a stable map with at most one cusp point such that the minimum $i_{0}$ of the absolute indices of folds satisfies $0<i_{0}<\lfloor(n-1)/2\rfloor$. Then, by using always-realizable moves we can make the fold image of absolute index $i_{0}$ outward-directed in such a way that the other folds have absolute index strictly greater than $i_{0}$, that their images are situated in the innermost region and that the resulting stable map has at most one cusp.
@@
Let $D\subset S^{2}$ be a $2$–disk containing $f(S(f))$. We can also take an annulus $A\cong[0,1]\times S^{1}$ in $D$ containing $f(S(f))$ in its interior. We use almost the same argument as in the proof of [3, Theorem 4.1] in order to deform the given stable map so that the fold image of absolute index $i_{0}$ is contained in $A$, and is outward-directed, and that the other fold images are situated in the inner region of $D\setminus A$. The points different from the situation in [3] that we need to take care of are as follows.
@@
Let $f:M\to S^{2}$ be a stable map with at most one cusp point such that the minimum $i_{0}$ of the absolute indices of folds satisfies $0<i_{0}<\lfloor(n-1)/2\rfloor$. We assume that $M$ is connected. Then, by using always-realizable moves we can eliminate the folds of absolute index $i_{0}$ in such a way that the other folds have absolute index strictly greater than $i_{0}$ and that the resulting stable map has at most one cusp.
@@
We further follow the procedure as described in [3, Figs. 30 and 31]. If the top strand, denoted by $b$, does not intersect the other strands, then we can isotope it into a small open disk neighborhood of the north pole $NP$. Then, we can use the move as described in Fig. 10 of the present paper so as to turn it to a fold image of absolute index $i_{0}+1$ which is an embedded small circle and is directed outward with respect to $NP$. If the top strand does intersect the others, then we perform the moves as described in [3, Fig. 33], by using the move as in Fig. 11 of the present paper instead of the move as described in [3, Fig. 34]. Note that by this move, we again create an embedded fold image circle of index $i_{0}+1$, which can be isotoped to a neighborhood of the south pole $SP\in R$. In order to reach the second stage of [3, Fig. 33], we need to swing the relevant fold image of index $i_{0}$ over the region around $SP$: this is possible, since all the fold images in this region have absolute index strictly greater than $i_{0}$.
Then, we reach the final stage of [3, Fig. 35]. Finally, we apply the move as in Fig. 10 of the present paper to turn the index $i_{0}$ circle to one of index $i_{0}+1$.
@@
First, let us consider the case $n\geq 6$. If we proceed as in the proof of Lemma 4.10, then there may appear some small embedded fold image circles of index $i_{0}+1$ normally oriented outward (see Figure 11). Note that if we reverse the normal orientation, then the index turns into $i_{0}$, the absolute index, since $n$ is even and $i_{0}=(n-2)/2$. We used push moves under the presence of Lefschetz critical points (see [3, Fig. 2]). In our present case, we can instead use the moves as depicted in Fig. 12.
@@
Then, we reach the situation as described in the leftmost figure of [3, Fig. 36], where the absolute indices are all equal to $i_{0}$, the singular point set image is topologically embedded, and each component is normally oriented inward. Note that one of the components may contain a cusp image.
This completes the proof for the case $n\geq 6$.
When $n=4$, we first follow the procedure as described in the proof of Lemma 4.8, taking into account the following points. Note that in this case $i_{0}=1$.
(1) If we have a cusp, then we can use cusp-fold crossing moves which increase the number of crossings. Note that they are always-realizable moves. In our process, these moves are sufficient for our purpose.
(2) In our process, we create fold circles of absolute index $i_{0}+1$ whose images are embedded and are outward-directed. If we reverse the normal orientations, then they turn to embedded fold circles of absolute index $i_{0}=1$, which are inward-directed. Then, we can safely use the moves as described in Fig. 12.
As a result, we get a stable map $f$ with at most one cusp point such that the fold image of absolute index $i_{0}$ is outward-directed except for some small embedded inward-directed circles of absolute index $i_{0}$. Then, when we follow the procedure as described in the proof of Lemma 4.10, we put the small embedded circles in the region $R$. Then we can show that the fibers over the points $R$ adjacent to the fold image normally directed toward $NP$ are connected.
Consequently, we can follow the same procedure as in the case $n\geq 6$. This completes the proof. ∎